Each set of rahul clothes occupues space of 1500 CM he want to carry 7set of clothes and his camera which is a cube of edge 20 CM will the bag be enough for him
1
Expert's answer
2019-12-25T17:16:29-0500

Find the volume of the camera


"V_{cube}=side\times side\times side"

"V_{cube}=20 cm\times 20cm\times 20cm=8000\ cm^3"

What volume do 7set of clothes occupy?


"V_{7clothes}=7\times 1500\ cm^3=1050\ cm^3"

Since "V_{cube}=8000\ cm^3 <1050\ cm^3=V_{7clothes}" , then the bag will not be enough for Rahul.